OpenStudy (anonymous):

anyone have some good hints when finding the second derivative, I always get lost :(

OpenStudy (anonymous):

With what question?

OpenStudy (anonymous):

\[\sqrt[3]{x^2-1}\]

OpenStudy (anonymous):

Well I would make that (x^2-1)^(1/3)

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }(x^2-1)^{-2/3}(2x)\]

OpenStudy (anonymous):

Then take the derivative combining the power rule and the chain rule.

OpenStudy (anonymous):

\[\frac{ 2x }{ 3\sqrt[3]{(x^2-1)^2} }\]

OpenStudy (anonymous):

Yeah that's right.

OpenStudy (anonymous):

\[\frac{ 3(x^2-1)^{-2/3}(2)-2x(-2(x^2-1)^{-4/3}(2x) }{ 3\sqrt[3]{(x^2-1)^2}^2 }\]

OpenStudy (anonymous):

*faints*

OpenStudy (anonymous):

Wait.

OpenStudy (anonymous):

Do you need the second derivative, or do you need just tips?

OpenStudy (anonymous):

well, i can look up the second derivative, im just wondering if there is an effective way to solve this, i always get lost

OpenStudy (anonymous):

Not exactly, It just comes with practice. Eventually when you get to calc iii, it starts becoming second nature. That's because for Diff. Eq., you must understand them well enough to work backwards. Sorry for no helpful tips.

zepdrix (zepdrix):

Consistent notation helps a little bit. Like I notice that in the bottom you're using a root symbol, and in the top you're using fractional exponents. I wouldn't mix those like that unless you're really comfortable with it. Personally I tend to use the check mark if its a square root, but write a fraction if it's any other power/root. I've seen you do a lot of these problems, you seem pretty great with your algebra and middle steps. I would just recommend getting really comfortable with the product and quotient rule.

OpenStudy (anonymous):

\[\frac{ 6 }{ \sqrt[3]{x^2-1} }+\frac{ 8x^2 }{ \sqrt[3]{(x^2-1)^5} }\]

OpenStudy (anonymous):

thats all over

zepdrix (zepdrix):

ew you gave them the check marks XD lol

OpenStudy (anonymous):

\[3\sqrt[3]{(x^2-1)^2}\]

OpenStudy (anonymous):

thanks zep, i will work on it more :)

OpenStudy (anonymous):

so, can i cross cancel the top portion?

zepdrix (zepdrix):

Might wanna back up a lil bit and check your work :c hmm

OpenStudy (anonymous):

checking

zepdrix (zepdrix):